### v3

@@ 2.5 Balanced panel and pooled estimation
$\left(\widehat{\beta},\widehat{\alpha}\right)$ $=\mathop{\mathrm{argmin}}_{\big{\{}\beta\in\mathbb{R}^{h},\,\alpha\in\mathbb{R}^{n\times(q+1)}\big{\}}}\,\sum_{i=1}^{n}\sum_{t=\tau-R+1}^{\tau+h}\left(y_{it}-\sum_{k=1}^{h}{1}\{t=\tau+k\}\,\beta_{k}-\sum_{k=0}^{q}\alpha_{ik}\,t^{k}\right)^{2},$ (19)
@@
The alternative estimation strategies in (18) and (19) provide algebraically identical treatment effect estimates in the case of our baseline setting with $R=R_{i}$ and $q=q_{i}$. In a more general setting, however, it is possible to show that these alternative estimation strategies do not give the same treatment effect estimator, and may indeed give inconsistent estimates for ATT if applied incorrectly.
3 Extension: no control group, covariates
In this section, we consider different ways of incorporating covariates, including lagged outcomes, in the estimation of FAT. Compared to the baseline case, this requires specifying a model for the counterfactuals and imposing additional assumptions.
3.1 Homogeneous coefficients
Consider the following model for the counterfactual $y_{it}(0)$:
@@ 3.1 Homogeneous coefficients
Assuming that we have estimates $\widehat{\beta}$ for the common parameter that are consistent as $n\rightarrow\infty$ under correct model specification [^26] For example, when $q_{i}=0$ and $x_{it}=(y_{it-1},z^{\prime}_{it})^{\prime}$, a consistent estimator for $\beta=(\rho,\theta^{\prime})^{\prime}$ can be obtained by applying an IV regression to the first-differenced model $y_{it}-y_{it-1}$ $=\left[y_{it-1}-y_{it-2}\right]\,\rho+\left[z_{it}-z_{it-1}\right]^{\prime}\theta+\varepsilon_{it}-\varepsilon_{it-1},$ using, for example, $y_{it-2}$ and $z_{it-1}$ as instruments. In the Monte Carlo simulations, we further extend this case to $q_{i}=1$. leads to the model-based FAT, defined as:
@@
$\widehat{y}_{h}(\widehat{\beta},y_{i},x_{i})$ $:=x_{i\tau+h}^{\prime}\,\widehat{\beta}+\sum_{k=0}^{q_{i}}(\tau+h)^{k}\,\widehat{c}_{ik}^{(q_{i}R_{i})}(\widehat{\beta}),$ (23)
$\widehat{c}_{i}^{(q_{i},R_{i})}(\widehat{\beta})$ $:=\mathop{\mathrm{argmin}}_{c\in\mathbb{R}^{q_{i}+1}}\sum_{t\in{\cal T}_{i}}\left(y_{it}-x_{it}^{\prime}\,\widehat{\beta}-\sum_{k=0}^{q_{i}}t^{k}\,c_{k}\right)^{2},$ (24)
@@
    The forecast is unbiased when evaluated at the true parameter value $\beta_{0}$, i.e.,
    $\mathbb{E}\left[\widehat{y}_{h}(\beta_{0},y_{i},x_{i})-y_{i\tau+h}\left(0\right)\right]=0.$
@@
    The function $\widehat{y}_{h}(\beta,y_{i},x_{i})$ is twice continuously differentiable such that $\frac{\partial\widehat{y}_{h}(\beta_{0},y_{i},x_{i})}{\partial\beta}$ has finite second moments, and for some $\delta>0$ we have
    $R_{n}:=\sup_{\left\{\beta\,:\,\left\|\beta-\beta_{0}\right\|\leq\delta\right\}}\left\|\frac{1}{n}\sum_{i=1}^{n}\frac{\partial^{2}\widehat{y}_{h}(\beta,y_{i},x_{i})}{\partial\beta\partial\beta^{\prime}}\right\|=o_{P}(n^{1/2}).$
@@
    satisfies a CLT in the sense that
    $\frac{\frac{1}{\sqrt{n}}\sum_{i}\left(u^{*}_{i\tau+h}-\mathbb{E}u^{*}_{i\tau+h}\right)}{\bar{\sigma}^{*}_{n}}\Rightarrow{\cal N}\left(0,1\right),\quad\bar{\sigma}^{*2}_{n}:={\rm Var}\left(\frac{1}{\sqrt{n}}\sum_{i}u^{*}_{i\tau+h}\right)<\infty.$
@@
3.2 Heterogeneous coefficients
If the model for the counterfactuals is an AR(p) with heterogeneous parameters, the time series literature (e.g., Fuller and Hasza (1980), Dufour (1984)), has derived conditions under which forecasts from an individual AR(p) model are unbiased. The assumptions are stationarity of the initial condition and symmetry of the error term.
@@ 3.2 Heterogeneous coefficients
4 Simulation study
Throughout this section, we set $b_{k}(t)=t^{k}$ in Definition 1, so that the forecast is:
$\widehat{y}_{i\tau+h}^{\rm PR}$ $:=\sum_{k=0}^{q_{i}}\widehat{c}_{ik}^{(q_{i},R_{i})}(\tau+h)^{k},$ $\widehat{c_{i}}^{(q_{i},R_{i})}:=\mathop{\mathrm{argmin}}_{c\in\mathbb{R}^{q_{i}+1}}\sum_{t\in{\cal T}_{i}}\left(y_{it}-\sum_{k=0}^{q_{i}}c_{k}\,t^{k}\right)^{2},$
where we suppressed the dependence on $(q_{i},R_{i})$ of $\widehat{y}_{i\tau+h}^{\rm PR}$. We refer to the associated estimator as the polynomial-regression FAT, defined as:
$\widehat{\rm FAT}_{h}^{\rm PR}:=\frac{1}{n}\sum_{i=1}^{n}\left(y_{i\tau+h}-\widehat{y}_{i\tau+h}^{\rm PR}\right).$ (29)
All simulations use the same tuning parameters $q$ and $R$ across individuals.
4.1 Polynomial-regression and model-based FAT with nonstationary initial condition
The goal of this simulation is twofold: to understand the impact of a nonstationary initial condition and to compare the performance of the polynomial-regression FAT in (29) to that of the model-based FAT in (22). The nonstationary initial condition violates Assumption 4, making the polynomial-regression FAT inconsistent, whereas we consider a DGP where it is possible to obtain a consistent model-based FAT using an appropriate estimator of the model’s parameters.
The DGP specifies the counterfactual for $i=1,\dots,n$ as:
$y_{it}(0)$ $=y_{it}^{(1)}(0)+y_{it}^{(3)}(0),\;t=1,\dots,T,$
$y_{it}^{(1)}(0)$ $=\mu_{i}+\rho y_{it-1}^{(1)}(0)+u_{it},\;t\geq 1,\;y_{i0}^{(1)}(0)\sim\mathcal{N}\left(1,2\right),$
$y_{it}^{(3)}(0)$ $=\delta_{i}t,$
$\mu_{i}$ $\sim\mathcal{U}\left[-1,1\right],\;u_{it}\sim\mathcal{N}\left(0,1\right),\;\delta_{i}=1,\;\rho\in\left\{0.2,0.9\right\},$
where $y_{it}^{(1)}(0)$ is an autoregressive process with the initial condition not drawn from the stationary distribution, and $y_{it}^{(3)}(0)$ is a deterministic linear time trend with homogeneous coefficients. [^28] Here, $\mu_{i}$ is not correlated with the initial condition. Additional Monte Carlo results, available upon request, show similar findings when the fixed effects are correlated with the initial condition. Here, $y_{it}(1)=y_{it}(0)$ at $t=\tau+1$, so ${\rm ATT}_{i\tau+1}=0$.
We consider a balanced panel with $T=6$ periods, with $\tau=5$, so that the first $5$ periods are the “pre-treatment,” and the last period is the “post-treatment” period. We focus on the $h=1$ case and show results for two sample sizes $n\in\left\{50,1000\right\}$.
The polynomial-regression FAT is computed for each $i$ by regressing $\left\{y_{it}\right\}_{t=1}^{5}$ on a polynomial in $t$ of order $q\leq T-2$, and then computing the forecast $\widehat{y}^{\rm PR}_{i6}\left(0\right)=\sum_{k=0}^{q}\widehat{c}_{ik}^{\left(q,q+1\right)}6^{k}$. We then use this individual forecast to compute $\widehat{\rm FAT}_{1}^{\rm PR}$ as in (29).
The model-based FAT is computed as described in Section 3.1, where the common AR parameter $\rho$ is estimated via Anderson-Hsiao, which allows for a nonstationary initial condition. We estimate $\rho$ in two different ways. We use $y_{it-3}$ as an instrument when the linear time trend is correctly accounted for, and $y_{it-2}$ as an instrument when the linear time trend is not accounted for. Given the estimate $\widehat{\rho}$, we then regress $\left\{y_{it}-\widehat{\rho}y_{it-1}\right\}_{t=1}^{5}$ on a polynomial in $t$ of order $q\leq T-2$ for each $i$, and then compute the forecast $\widehat{y}^{\rm MB}_{i6}\left(0\right)=\widehat{\rho}y_{i5}+\sum_{k=0}^{q}\widehat{c}_{ik}^{\left(q,q+1\right)}6^{k}$. $\widehat{\rm FAT}_{1}^{\rm MB}$ is then computed as in (22).
Here, the polynomial-regression FAT is inconsistent, while the model-based FAT is consistent when the correct instrument is used in the estimation of $\rho$ in the first step, but it is inconsistent when the wrong instrument is used in the first step.
Table 1 shows the bias and the standard error (computed across the simulations) of the polynomial-regression FAT (PR) and of the model-based FAT when using the correct instrument (MB) and when using the wrong instrument (MB missp.) in the first step. We see that the model-based FAT is sensitive to misspecification, with potentially very large biases and standard errors when using an inconsistent estimator in the first step. Even when using the correct instrument, we see that there can still be bias due to estimation error in $\widehat{\rho}$ (the so-called Nickell bias). [^29] In the literature on dynamic panel models, it is known that the Anderson-Hsiao estimator suffers from finite-sample bias when $\rho$ is close to unity and $T$ is small. One could try to address the latter issue via a bias correction method (e.g. Juodis (2013)). However, to our knowledge, none of the bias correction methods for short-$T$ allow for unit root components and they all maintain correct model specification. On the other hand, the polynomial-regression FAT has good performance for $q=2,3$, in spite of violation of its underlying assumptions. This suggests that the polynomial-regression FAT is robust and that higher order polynomials could guard against nonstationary initial conditions. Arguably, by considering an estimator of $\rho$ in a linear dynamic model with short-$T$, this simulation study provides a reminder of both the difficulties and the importance of finding dynamic short-panel estimators with good properties. Perhaps the lesson to be drawn is that the Nickell bias in $\widehat{\rho}$ shows up as bias in $\widehat{{\rm FAT}}^{MB}$, while an added advantage of $\widehat{{\rm FAT}}^{PR}$ is that it by-passes the incidental parameter problem by not relying on a first-stage estimate of the model’s parameters.
$\rho=0.2$ $q=0$ $q=1$ $q=2$ $q=3$
$n=50$ PR
$1.25$
$(0.19)$
$0.02$
$(0.31)$
$0.003$
$(0.54)$
$0.01$
$(0.99)$
MB
$1.25$
$(1.26)$
$0.01$
$(0.37)$
$0.02$
$(0.54)$
$0.03$
$(0.89)$
MB
missp.
$41.95$
$(1292.09)$
$-0.86$
$(22.45)$
$-0.14$
$(2.13)$
$-0.47$
$(6.77)$
$n=1000$ PR
$1.25$
$(0.04)$
$0.001$
$(0.07)$
00
$(0.12)$
$0.01$
$(0.24)$
MB
$1.00$
$(0.19)$
$0.001$
$(0.08)$
$-0.01$
$(0.14)$
$0.003$
$(0.27)$
MB
missp.
$0.23$
$(0.1)$
$-0.003$
$(0.11)$
00
$(0.21)$
$-0.02$
$(0.41)$
$\rho=0.9$
$n=50$ PR
$4.69$
$(0.16)$
$0.61$
$(0.21)$
$-0.06$
$(0.36)$
$0.01$
$(0.65)$
MB
$1.03$
$(3.38)$
$-0.04$
$(0.69)$
$-0.17$
$(0.75)$
$-0.25$
$(1.42)$
MB
missp.
$186.7$
$(5905.57)$
$2.36$
$(86.1)$
$0.2$
$(10.98)$
$-0.42$
$(15.9)$
$n=1000$ PR
$4.69$
$(0.03)$
$0.59$
$(0.05)$
$-0.06$
$(0.08)$
$0.01$
$(0.15)$
MB
$1.03$
$(0.43)$
$-0.003$
$(0.14)$
$-0.01$
$(0.14)$
$-0.01$
$(0.28)$
MB
missp.
$0.96$
$(45.95)$
$-4.58$
$(157.56)$
$0.12$
$(2.27)$
$-4.21$
$(135.57)$
Table 1: Bias and standard error (in parentheses) for the polynomial-regression FAT (PR), the model-based FAT that takes account of the linear time trend and uses the correct instruments in the first step (MB), and the model-based FAT that uses the incorrect instrument in the first step (MB missp.). $n$ is the sample size and $q$ is the polynomial order.
4.2 Choice of tuning parameters for polynomial-regression FAT
In this section, we compare the finite-sample performance of the polynomial-regression FAT estimator across different tuning parameters: the polynomial order, $q$, and the estimation window, $R$, for different specifications of the DGP. All specifications satisfy Assumption 4, with the counterfactual process specified as the sum of up to three different components.

### v4

@@ 2.5 Balanced panels and pooled estimation
[…]
@@
3 Parametric-Model FAT
In this section, we consider an alternative approach that explicitly incorporates covariates, including lagged outcomes, in the estimation of FAT. Compared to the baseline case, this involves specifying a parametric model for the counterfactuals and imposing additional assumptions.
3.1 Homogeneous parameters
Consider the following parametric model for the counterfactual $y_{it}(0)$:
@@ 3.1 Homogeneous parameters
Assuming that we have estimates $\widehat{\beta}$ for the common parameter that are consistent as $n\rightarrow\infty$ under correct model specification [^13] For example, when $q_{i}=0$ and $x_{it}=(y_{it-1},z^{\prime}_{it})^{\prime}$, a consistent estimator for $\beta=(\rho,\theta^{\prime})^{\prime}$ can be obtained by applying an IV regression to the first-differenced model $y_{it}-y_{it-1}$ $=\left[y_{it-1}-y_{it-2}\right]\,\rho+\left[z_{it}-z_{it-1}\right]^{\prime}\theta+\varepsilon_{it}-\varepsilon_{it-1},$ using, for example, $y_{it-2}$ and $z_{it-1}$ as instruments. In the Monte Carlo simulations, we further extend this case to $q_{i}=1$. leads to the parametric-model FAT, defined as:
@@
[…]
@@
    The forecast is unbiased when evaluated at the true parameter value $\beta_{0}$, i.e., E[^y_h(β_0,y_i,x_i) - y_iτ+h(0) ] = 0.
@@
[…]
@@
    satisfies a CLT in the sense that 1ni(u*iτ+h-Eu*iτ+h)¯σ*nN(0,1),  ¯σ^*2_n := Var(1n_i u^*_iτ+h )<.
@@
When the covariate vector includes lagged outcomes (e.g., $x_{it}=y_{i,t-1}$), the forecast in (23) requires these lagged values to be observed. For $h=1$, this poses no problem since $y_{i\tau}$ is observed. For $h\geq 2$, however, the required intermediate counterfactuals $y_{i,\tau+1}(0),\ldots,y_{i,\tau+h-1}(0)$ are not observed under treatment, so iterated forecasting is not directly feasible. One solution is to adopt a local-projection approach that directly models the $h$-step-ahead relationship between $y_{i,\tau+h}(0)$ and $x_{i\tau}$, analogous to Jordà (2005). Alternatively, one may use the Unobserved-Components FAT from Section 2, which does not require dynamic covariates and remains feasible at all horizons.
3.2 Heterogeneous parameters
If the model for the counterfactuals is an AR(p) with heterogeneous parameters, the time series literature (e.g., Fuller and Hasza 1980; Dufour 1984) has derived conditions under which forecasts from an individual AR(p) model are unbiased. The assumptions are stationarity of the initial condition and symmetry of the error term.
@@ 3.2 Heterogeneous parameters
4 Simulation studies
In this section, we study the finite-sample performance of our estimators. Although we focus on a universal-treatment setting, the insights from the simulations carry over unchanged to staggered adoption, as linearity of the forecasting operator ensures unbiased individual forecasts aggregate to unbiased cohort forecasts.
In Section 4.1, we compare the Unobserved-Components FAT (UC FAT) to the Parametric-Model FAT (PM FAT) across forecast horizons, sample sizes, autoregressive persistence, and alternative initial-condition regimes (stationary vs. nonstationary). PM FAT can inherit small-sample biases from first-stage estimation of the AR parameter, whereas UC FAT avoids these by construction; the contrast is most pronounced under high persistence and small $n$. When initial conditions are stationary, finite-sample biases are generally small across forecast horizons.
Section 4.2 focuses on UC FAT and evaluates how tuning parameter choices - polynomial order $q$ and pre-treatment window $R$ - trade off bias and variance under DGPs satisfying Assumptions 3 and 4. In the small-$T$ setting we consider, the simulations support the recommendation to use the lowest polynomial order that accommodates any deterministic trend in the outcome and, conditional on that choice, select the largest feasible pre-treatment window. When the trend order is underfit, bias emerges; in that case, shortening the pre-treatment window can mitigate misspecification. Once the polynomial order is rich enough, bias is largely insensitive to $(q,R)$, and precision improves with lower order and longer windows. These patterns are robust across DGPs with heterogeneous autoregressive parameters and individual-specific trends, and they provide clear guidance for empirical implementation.
4.1 Performance across forecast horizons
We consider the following DGP for the counterfactual outcome:
$y_{it}(0)\;=\;y^{(1)}_{it}(0)+y^{(3)}_{it}(0),\qquad t=1,\dots,T,$
where the autoregressive component has a homogeneous AR parameter
$y^{(1)}_{it}(0)=\mu_{i}+\rho\,y^{(1)}_{i,t-1}(0)+u_{it},\quad\mu_{i}\sim\mathcal{U}[-1,1],\quad u_{it}\sim\mathcal{N}(0,1),\quad\rho\in\{0.2,0.9\},$
and the deterministic trend is homogeneous linear
$y^{(3)}_{it}(0)=\delta_{i}t,\qquad\delta_{i}=1.$
We consider two regimes for $y^{(1)}_{i0}(0)$:
$y^{(1)}_{i0}(0)$ $\sim\mathcal{N}(1,2)$ $\text{(nonstationary)},$
$y^{(1)}_{i0}(0)$ $\sim\mathcal{N}\!\big(\tfrac{\mu_{i}}{1-\rho},\,\tfrac{1}{1-\rho^{2}}\big)$ $\text{(stationary)}.$
We work with a balanced panel, $T=8$ and $\tau=5$ (five pre-treatment periods), forecast horizons $h\in\{1,2,3\}$, and sample sizes $n\in\{50,1000\}$. Treatment effects at $\tau+h$ are zero by construction.
Estimators.
For $b_{k}(t)=t^{k}$, the UC forecast for unit $i$ at $\tau+h$ is
$\widehat{y}^{\text{UC}}_{i,\tau+h}(0)=\sum_{k=0}^{q}\widehat{c}_{ik}(\tau+h)^{k},\quad\widehat{c}_{i}=\arg\min_{c\in\mathbb{R}^{q+1}}\sum_{t=1}^{\tau}\Big(y_{it}-\sum_{k=0}^{q}c_{k}t^{k}\Big)^{2},$
and our UC FAT estimator is
$\widehat{\text{FAT}}^{\text{UC}}_{h}=\frac{1}{n}\sum_{i=1}^{n}\big(y_{i,\tau+h}-\widehat{y}^{\text{UC}}_{i,\tau+h}(0)\big).$ (29)
We compute (29) using the fatEstimator package in R, with $R=q+1$ and $q\in\{0,1,2,3\}$.
The PM forecast first estimates $\rho$ by Anderson–Hsiao using $y_{i,t-2}$ as an instrument for $\Delta y_{i,t-1}$ in $\Delta y_{it}=\rho\,\Delta y_{i,t-1}+\Delta u_{it}$, then fits a polynomial to the residuals $y_{it}-\widehat{\rho}\,y_{i,t-1}$, so that:
$\widehat{y}^{\text{PM}}_{i,\tau+h}(0)=\widehat{\rho}\,y_{i,\tau}+\sum_{k=0}^{q}\widehat{c}_{ik}(\tau+h)^{k},\quad\widehat{c}_{i}=\arg\min_{c}\sum_{t=1}^{\tau}\Big(y_{it}-\widehat{\rho}\,y_{i,t-1}-\sum_{k=0}^{q}c_{k}t^{k}\Big)^{2},$
with the PM FAT estimator given by:
$\widehat{\text{FAT}}^{\text{PM}}_{h}=\frac{1}{n}\sum_{i=1}^{n}\big(y_{i,\tau+h}-\widehat{y}^{\text{PM}}_{i,\tau+h}(0)\big).$ (30)
We estimate $\rho$ using the ivreg package in R and then use the fatEstimator package on the residuals $y_{it}-\widehat{\rho}\,y_{i,t-1}$. The specification fits polynomials of degree $q\in\{0,1,2,3\}$ to $R=q+1$ pre-treatment periods.
Simulation scenarios.
We consider four scenarios, varying the cross-sectional size and the autoregressive parameter:
$(n,\rho)\in\{(50,0.2),\ (50,0.9),\ (1000,0.2),\ (1000,0.9)\}.$
We report bias and RMSE for UC and PM FAT across polynomial orders $q\in\{0,1,2,3\}$, forecast horizons $h\in\{1,2,3\}$, AR parameter magnitude $\rho\in\{0.2,0.9\}$, and sample size $n\in\{50,1000\}$, based on 500 Monte Carlo replications.
Tables 1 and 2 report bias and RMSE for the UC FAT for a DGP with initial condition drawn from the nonstationary distribution (Table 1) and from the stationary distribution (Table 2). With a nonstationary initial condition (Table 1), the $q=0$ specification - which omits the linear trend - exhibits large upward bias that grows roughly linearly with the forecast horizon $h$. Including a linear term ($q=1$) effectively removes this bias: the remaining deviations from zero are small across all horizons and values of $\rho$. Higher-order specifications ($q=2,3$) do not reduce asymptotic bias in this DGP (the true trend is linear), but they affect small-sample behavior: they can attenuate some initialization effects at the cost of substantially larger RMSE because of polynomial extrapolation. In all cases, RMSE increases with the forecast horizon.
With a stationary initial condition (Table 2), the $q=0$ specification still omits the linear time trend and exhibits a systematic upward bias. Once a linear term is included ($q=1$), the forecast specification matches the deterministic component of the DGP, and bias is essentially zero across horizons and values of $\rho$; the small nonzero entries in the table reflect finite-sample variability. Higher-order specifications ($q=2,3$) remain asymptotically unbiased but exhibit larger RMSE due to higher-order polynomial extrapolation. As expected, RMSE increases with $h$, while bias patterns are stable across $\rho$.
Tables 3 and 4 report bias and RMSE of PM FAT for a DGP with initial condition drawn from the nonstationary distribution (Table 3) and from the stationary distribution (Table 4). With a nonstationary initial condition (Table 3), the $q=0$ specification again omits the linear trend and generates a large upward bias that increases with the forecast horizon $h$. Introducing a linear term ($q=1$) removes this trend-misspecification bias at the population level, so that any remaining bias is driven by first-stage estimation of the AR parameter $\rho$. In finite samples, PM FAT inherits a small Nickell-type bias through $\hat{\rho}$, which propagates into the $h$-step forecast and yields modest biases even when $q=1$. Higher-order specifications ($q=2,3$) do not further reduce this asymptotic bias but amplify RMSE because polynomial extrapolation increases forecast variance, especially at longer horizons. As with UC FAT, RMSE rises with $h$, while the qualitative bias pattern is similar across values of $\rho$.
With a stationary initial condition (Table 4), the $q=0$ model still omits the linear trend and exhibits a systematic upward bias, though somewhat attenuated relative to the nonstationary case. Once a linear term is included ($q=1$), the forecasting specification matches the deterministic component of the DGP, and PM FAT is asymptotically unbiased; the small residual biases reflect the Nickell-type distortion from estimating $\rho$ in the first step. As before, higher-order specifications ($q=2,3$) remain asymptotically unbiased but exhibit larger RMSE due to variance inflation from higher-order polynomials. Compared to the nonstationary case, the stationary initial condition reduces the magnitude of the finite-sample biases, while RMSE continues to increase with $h$ and bias patterns remain stable across $\rho$.
Table 1: Bias and RMSE for Unobserved-Components FAT across forecast horizons and polynomial orders using $R=q+1$ pre-treatment periods. The DGP features a nonstationary initial condition and a homogeneous AR parameter.
$q=0$ $q=1$ $q=2$ $q=3$
$n$ $\rho$ $h$ Bias RMSE Bias RMSE Bias RMSE Bias RMSE
50 0.2 1 $2.0036$ $2.0133$ $0.0165$ $0.4645$ 0.0746-0.0746 $1.2652$ $0.3312$ $3.3415$
2 $2.9911$ $2.9991$ $0.0066$ $0.6930$ 0.1439-0.1439 $2.4915$ $0.9446$ $8.0830$
3 $3.9955$ $4.0011$ 0.0090-0.0090 $0.8508$ 0.4136-0.4136 $3.8825$ $1.0853$ $14.6870$
50 0.9 1 $1.8777$ $1.8906$ $0.0316$ $0.3525$ $0.0029$ $0.8469$ 0.0107-0.0107 $2.1449$
2 $2.8088$ $2.8229$ $0.0338$ $0.5405$ 0.0221-0.0221 $1.6945$ $0.0071$ $5.2861$
3 $3.7749$ $3.7879$ $0.0435$ $0.6605$ 0.1310-0.1310 $2.6204$ 0.4783-0.4783 $9.7770$
1000 0.2 1 $1.9964$ $1.9969$ $0.0029$ $0.1056$ 0.0901-0.0901 $0.2949$ $0.2897$ $0.7752$
2 $2.9996$ $2.9999$ $0.0228$ $0.1512$ 0.1186-0.1186 $0.5365$ $0.9476$ $1.8948$
3 $3.9981$ $3.9984$ $0.0138$ $0.1849$ 0.2710-0.2710 $0.8548$ $1.7317$ $3.5273$
1000 0.9 1 $1.8730$ $1.8736$ $0.0149$ $0.0808$ 0.0207-0.0207 $0.1926$ 0.0412-0.0412 $0.4738$
2 $2.8187$ $2.8194$ $0.0443$ $0.1229$ $0.0039$ $0.3625$ $0.0390$ $1.0985$
3 $3.7728$ $3.7735$ $0.0635$ $0.1613$ 0.0456-0.0456 $0.5631$ 0.0826-0.0826 $2.0449$
  • Results are based on $500$ Monte Carlo replications of the DGP described in Section 4.1. For $q=0$, bias grows linearly in $h$ and is large. Once a linear trend is included ($q=1$), bias is substantially reduced and remains small across horizons. Higher-order specifications ($q=2,3$) do not remove additional bias in this DGP but inflate forecast variability.
Table 2: Bias and RMSE for Unobserved-Components FAT across forecast horizons and polynomial orders using $R=q+1$ pre-treatment periods. Stationary initial condition and homogeneous AR parameter.
$q=0$ $q=1$ $q=2$ $q=3$
$n$ $\rho$ $h$ Bias RMSE Bias RMSE Bias RMSE Bias RMSE
50 0.2 1 $1.9829$ $1.9928$ 0.0598-0.0598 $0.4593$ 0.1349-0.1349 $1.2571$ 0.2328-0.2328 $3.2557$
2 $2.9943$ $3.0014$ 0.0196-0.0196 $0.6481$ 0.0366-0.0366 $2.2848$ 0.0147-0.0147 $7.3752$
3 $4.0111$ $4.0164$ $0.0473$ $0.8165$ $0.1694$ $3.6622$ $0.3981$ $14.2240$
50 0.9 1 $1.9962$ $2.0072$ 0.0394-0.0394 $0.3575$ 0.1069-0.1069 $0.8477$ 0.2077-0.2077 $2.1326$
2 $2.9950$ $3.0050$ 0.0163-0.0163 $0.5070$ 0.0422-0.0422 $1.5812$ 0.0577-0.0577 $4.8478$
3 $4.0120$ $4.0218$ $0.0422$ $0.6725$ $0.1414$ $2.5028$ $0.3873$ $9.3314$
1000 0.2 1 $2.0005$ $2.0010$ 0.0002-0.0002 $0.1117$ 0.0032-0.0032 $0.2971$ 0.0147-0.0147 $0.7600$
2 $2.9976$ $2.9979$ 0.0038-0.0038 $0.1482$ 0.0216-0.0216 $0.5316$ 0.0883-0.0883 $1.6581$
3 $3.9998$ $4.0000$ 0.0127-0.0127 $0.1847$ 0.0576-0.0576 $0.8121$ 0.2343-0.2343 $3.0834$
1000 0.9 1 $2.0013$ $2.0017$ $0.0008$ $0.0832$ 0.0007-0.0007 $0.2051$ 0.0066-0.0066 $0.5019$
2 $2.9980$ $2.9985$ 0.0029-0.0029 $0.1125$ 0.0091-0.0091 $0.3661$ 0.0492-0.0492 $1.1188$
3 $4.0034$ $4.0039$ 0.0073-0.0073 $0.1460$ 0.0420-0.0420 $0.5682$ 0.1535-0.1535 $2.0527$
  • Results are based on $500$ Monte Carlo replications of the DGP described in Section 4.1. For $q=0$, the estimator omits the linear time trend and produces a systematic upward bias. Once a linear term is included ($q=1$), the forecast specification matches the deterministic component of the DGP, so bias vanishes; small deviations in the table reflect finite-sample variability. Higher-order specifications ($q=2,3$) remain unbiased asymptotically in $n$ but yield larger RMSE because polynomial extrapolation amplifies forecast variance.
Table 3: Bias and RMSE for Parametric-Model FAT across forecast horizons and polynomial orders using $R=q+1$ pre-treatment periods. Nonstationary initial condition and homogeneous AR parameter.
$q=0$ $q=1$ $q=2$ $q=3$
$n$ $\rho$ $h$ Bias RMSE Bias RMSE Bias RMSE Bias RMSE
50 0.2 1 $1.5740$ $1.6295$ $0.0037$ $0.5433$ 0.0259-0.0259 $1.5487$ 0.1559-0.1559 $4.0561$
2 $2.3543$ $2.4345$ 0.0251-0.0251 $0.7298$ 0.2214-0.2214 $2.7906$ 0.8179-0.8179 $8.9927$
3 $3.1187$ $3.2246$ $0.0497$ $0.9475$ $0.1381$ $4.1904$ 0.1103-0.1103 $15.8752$
50 0.9 1 $0.8352$ $1.4225$ 0.0538-0.0538 $0.5179$ 0.0828-0.0828 $1.4012$ 0.1100-0.1100 $3.5624$
2 $1.3678$ $2.1894$ 0.0508-0.0508 $0.6659$ 0.2010-0.2010 $2.4168$ 0.5022-0.5022 $7.7318$
3 $1.7103$ $2.8662$ 0.0006-0.0006 $0.8940$ $0.1780$ $3.7222$ $0.8174$ $13.8606$
1000 0.2 1 $1.5999$ $1.6028$ $0.0107$ $0.1181$ $0.0264$ $0.3309$ $0.0428$ $0.8568$
2 $2.4025$ $2.4061$ $0.0012$ $0.1462$ 0.0091-0.0091 $0.5638$ 0.0176-0.0176 $1.9492$
3 $3.1996$ $3.2046$ $0.0028$ $0.2096$ $0.0177$ $0.9540$ $0.0418$ $3.5432$
1000 0.9 1 $0.2371$ $0.3189$ $0.0084$ $0.1166$ $0.0243$ $0.3281$ $0.0519$ $0.8504$
2 $0.3438$ $0.4728$ 0.0018-0.0018 $0.1464$ 0.0059-0.0059 $0.5634$ $0.0306$ $1.9445$
3 $0.4719$ $0.6387$ 0.0014-0.0014 $0.2074$ $0.0171$ $0.9476$ $0.1110$ $3.5020$
  • Results are based on $500$ Monte Carlo replications of the DGP described in Section 4.1. For $q=0$, the estimator omits the linear time trend and produces a systematic upward bias. Once a linear term is included ($q=1$), the bias disappears. Higher-order specifications ($q=2,3$) remain unbiased asymptotically in $n$ but yield larger RMSE. RMSE grows with $h$, while bias patterns remain stable across different values of $\rho$.
Table 4: Monte Carlo Results: Bias and RMSE for Parametric-Model FAT across forecast horizons and polynomial using $R=q+1$ pre-treatment periods. Stationary initial condition and homogeneous AR parameter.
$q=0$ $q=1$ $q=2$ $q=3$
$n$ $\rho$ $h$ Bias RMSE Bias RMSE Bias RMSE Bias RMSE
50 0.2 1 $1.6119$ $1.6686$ $0.0154$ $0.5343$ 0.0164-0.0164 $1.4927$ 0.1151-0.1151 $3.9297$
2 $2.4122$ $2.4798$ $0.0148$ $0.7293$ 0.1584-0.1584 $2.6326$ 0.7999-0.7999 $8.4156$
3 $3.1723$ $3.2695$ $0.0709$ $0.9797$ $0.2025$ $4.5242$ $0.1281$ $17.0235$
50 0.9 1 $0.7370$ $1.2933$ $0.0108$ $0.4967$ $0.0084$ $1.3603$ $0.0136$ $3.5382$
2 $1.1573$ $1.9940$ $0.0348$ $0.6901$ 0.0844-0.0844 $2.4447$ 0.5345-0.5345 $7.7359$
3 $1.6071$ $2.7400$ $0.1014$ $0.9006$ $0.3299$ $4.0615$ $0.7651$ $15.2754$
1000 0.2 1 $1.6071$ $1.6097$ $0.0142$ $0.1143$ $0.0309$ $0.3139$ $0.0634$ $0.8297$
2 $2.3981$ $2.4016$ $0.0087$ $0.1664$ $0.0206$ $0.5888$ $0.0087$ $1.9239$
3 $3.1944$ $3.1981$ 0.0003-0.0003 $0.2023$ $0.0102$ $0.9322$ $0.0347$ $3.6121$
1000 0.9 1 $0.2328$ $0.3175$ $0.0145$ $0.1136$ $0.0327$ $0.3122$ $0.0680$ $0.8272$
2 $0.3432$ $0.4814$ $0.0090$ $0.1653$ $0.0231$ $0.5867$ $0.0129$ $1.9199$
3 $0.4503$ $0.6249$ $0.0014$ $0.2016$ $0.0165$ $0.9265$ $0.0547$ $3.5848$
  • With stationary initial conditions, the $q=0$ specification still produces upward bias from omitting the linear trend. Once a linear term is included ($q=1$), the PM estimator is asymptotically unbiased; the small residual biases reflect estimation of $\rho$ in the first step. Higher-order $q$ remain unbiased but exhibit larger RMSE due to variance inflation. Compared to the nonstationary case, the stationary initial condition attenuates the size of the finite-sample biases.
4.2 Choice of tuning parameters for Unobserved-Components FAT
In this section, we compare the finite-sample performance of the Unobserved-Components FAT estimator across different tuning parameters: the polynomial order, $q$, and the estimation window, $R$, for different specifications of the DGP. All specifications satisfy Assumption 4, with the counterfactual process specified as the sum of up to three different components.